### v1

# In-Context Learning Dynamics with Random Binary Sequences

Abstract
Large language models (LLMs) trained on huge corpora of text datasets demonstrate complex, emergent capabilities, achieving state-of-the-art performance on tasks they were not explicitly trained for. The precise nature of LLM capabilities is often unclear, and different prompts can elicit different capabilities through in-context learning. We propose a Cognitive Interpretability framework that enables us to analyze in-context learning dynamics to understand latent concepts in LLMs underlying behavioral patterns. This provides a more nuanced understanding than success-or-failure evaluation benchmarks, but does not require observing internal activations as a mechanistic interpretation of circuits would require. Inspired by the cognitive science of human randomness perception, we use random binary sequences as context and study dynamics of in-context learning by manipulating properties of context data, such as sequence length. In the latest GPT-3.5+ models, we find emergent abilities to generate pseudo-random numbers and learn basic formal languages, with striking in-context learning dynamics where model outputs transition sharply from pseudo-random behaviors to deterministic repetition.
1 Introduction
Large language models (LLMs) achieve high performance on a wide variety of tasks they were not explicitly trained for, demonstrating complex, emergent capabilities (Brown et al., 2020; Radford et al., 2019; Wei et al., 2022a; 2023; b; Lu et al., 2023; Bommasani et al., 2021; Wei et al., 2021; Kojima et al., 2022; Pan et al., 2023; Pallagani et al., 2023; Li et al., 2022). In-context learning (ICL) describes how task-specific patterns of behavior are elicited in LLMs by different prompts (or contexts(Brown et al., 2020; Wei et al., 2022b; Dong et al., 2022; Zhou et al., 2022; Wang et al., 2023; Shin et al., 2022; Min et al., 2022; Si et al., 2023; Xie et al., 2022; Zhang et al., 2023b). Although no weight updates occur in ICL, different input contexts can activate, or re-weight, different latent algorithms in an LLM, analogous to how traditional learning methods such as gradient descent use training data to re-weight model parameters to learn representations (Goldblum et al., 2023; Dai et al., 2023; Ahn et al., 2023; Von Oswald et al., 2023; Akyürek et al., 2022; Li et al., 2023a). Two seemingly equivalent prompts can, however, evoke very different behaviors in LLMs (Si et al., 2023). ICL makes interpreting what representations in an LLM are being evoked by a given input more challenging than interpreting representations in task-specific language models, where the space of capabilities is more restricted and well-defined. Our central motivation here is to interpret emergent capabilities and latent concepts underlying complex behaviors in LLMs by analyzing in-context learning behavioral dynamics, without directly observing hidden unit activations or re-training models on varied datasets.
Inspired by computational approaches to human cognition (Tenenbaum, 1998; Griffiths & Tenenbaum, 2003; Piantadosi et al., 2012; Spivey et al., 2009), we model and interpret latent concepts evoked in LLMs by different contexts, without observing or probing model internals. This approach, which we call Cognitive Interpretability (see Fig. 1), is a middle ground between shallow test-set evaluation benchmarks on one hand (Srivastava et al., 2022; Saparov & He, 2022; Min et al., 2022; Lanham et al., 2023; Ganguli et al., 2023; Bowman et al., 2022; Kadavath et al., 2022; Perez et al., 2022; Prystawski & Goodman, 2023; Turpin et al., 2023) and mechanistic neuron- and circuit-level understanding of pre-trained models’ capabilities on the other (Nanda et al., 2023; Goh et al., 2021; Geva et al., 2020; Belinkov, 2022; Li et al., 2022; Wang et al., 2022; Chughtai et al., 2023; Gurnee et al., 2023; Foote et al., 2023; Lubana et al., 2023; Lieberum et al., 2023; Barak et al., 2022; Liu et al., 2022b; Zhong et al., 2023). The reliance on benchmarks loosely parallels early behaviorist psychology, when theories of human and animal learning only assumed stimulus-response associations, without positing theories about mental processes or neural substrates (Gardner, 1987). Circuit-level mechanistic interpretations of neural networks parallel neurobiology, which offers physical models of information processing in biological neurons, but typically does not account for the structure of high-level behavior. Cognitive interpretability, like cognitive science, is aimed at predicting behavioral outcomes over a potentially infinite space of possible tasks. It defines high-level specifications of behaviors performed by LLMs, which, we argue, should be the first step before mechanistic understanding of circuits existent in a model is pursued.
Here, we focus on cognitive interpretability of LLMs in the domain of random sequences of binary values. Random binary sequences are a minimal domain that have been studied extensively in statistics, formal language theory, and algorithmic information theory (Sipser, 1996; Chaitin, 1966; Li et al., 1997). We can use this domain to systematically test few-shot learning as a function of context length $|x|$ by testing different input sequences $x$. We can also test zero-shot learning by having models generate sequences with no context ($|x|=0$), without relying on alternate prompt formats such as chain-of-thought reasoning (Kojima et al., 2022; Wei et al., 2022b). Moreover, language generation trajectories over binary sequences can also be analyzed and visualized much more easily than typical user-chatbot interaction trajectories (Zhang et al., 2023a; Berglund et al., 2023), since the token-by-token branching factor is only two. Random binary sequences have also been a target domain in cognitive science (specifically, subjective randomness), where researchers have studied the mechanisms and concepts that underlie how people generate random binary sequences or evaluate the randomness of given sequences (Falk & Konold, 1997; Tversky & Gilovich, 1989; Griffiths et al., 2018).
Figure 1: Cognitive Interpretability in the context of prior work Cognitive Interpretability studies in-context learning dynamics in LLMs, positing theories of the latent concepts enabling behavioral capabilities. It is a middle ground between behavioral benchmarks, which treat models as black boxes and evaluate hit-or-miss accuracy, and mechanistic interpretability, which studies toy models and training loss dynamics, analogous to how cognitive science is a middle ground between behaviorist psychology and neurobiology in the study of human intelligence.
Computational cognitive scientists have related algorithmic information theory to human cognition, where mental concepts are viewed as programs, and cognitive hypothesis search over concepts is viewed as Bayesian inference (or Bayesian program learning) (Chater & Vitányi, 2003; Dehaene et al., 2022; Piantadosi et al., 2012; Goodman et al., 2008; Ullman et al., 2012; Ullman & Tenenbaum, 2020). In this vein, Griffiths & Tenenbaum (2003) and Griffiths et al. (2018) model subjective randomness in human cognition as probabilistic program induction, where a person must search over a space of non-random programs in order to answer the question, “was this sequence generated by a random process?” We argue that ICL can similarly be seen as under-specified program induction, where there is no single “correct” answer; instead, an LLM should appropriately re-weight latent algorithms. The domain of random sequences reflects this framing, in contrast to other behavioral evaluation methodologies, in that there is no correct answer to a random number generation or judgment task (Fig. 3). If the correct behavior is to match a target random process, then the right way to respond to a prompt Generate N flips from a fair coin is at best a uniform distribution over the tokens Heads and Tails, instead of a specific ground truth token sequence. If instead the correct behavior is to match how humans generate random sequences, the LLM must represent a more complex algorithm that matches the algorithm presumably being implemented by people. Overall, we make the following contributions in this work:
  • Evaluating ICL as Bayesian model selection in a practical setting. We find evidence of in-context learning operating as Bayesian model selection in LLMs in the wild, compared to prior work that trains smaller toy neural networks from scratch.
  • Subjective Randomness as a domain for studying ICL dynamics. We introduce a new domain, subjective randomness, which is both rich enough to be theoretically interesting, but simple enough that complex behavioral patterns are easily analyzed.
  • Sharp transitions in abilities to generate pseudo random numbers. We systematically analyze in-context learning dynamics in LLMs without observing or updating model weights, demonstrating sharp phase-changes in model behavior. This is a minimal, interpretable example of how the largest and most heavily fine-tuned LLMs can suddenly shift from one pattern of behavior to another during text generation.
2 Preliminaries: Bayesian Inference and In-Context Learning
Figure 2: Overview of our modeling framework. Given a pre-trained Large Language Model, we systematically vary input context prompts, in this case $x=(001)^{n}$ with $n=\{2,4\}$ (001001, $001$). The LM outputs $y$ (red) vary significantly as a function of $x$ (blue), based on some unknown latent concept space embedded in the LLM. We model a subset of the algorithms embedded in the LLM’s latent concept space $\mathcal{H}$ that are invoked by ICL (red + shaded box), to predict the LLM outputs $y$. With very little context ($x=\texttt{0}$), GPT-3.5+ generate subjectively random sequences, whereas with adequate context matching a simple formal language ($x=\texttt{001001001001}$), behavior follows the same deterministic pattern.
A multi-layer neural networks may possess multiple distributed circuits implementing computational primitives, such as basic mathematical operations like addition and sequence copying (Olsson et al., 2022; Nanda et al., 2023; Hanna et al., 2023; Zhong et al., 2023; Chan et al., 2022; Bietti et al., 2023). In state of the art LLMs with hundreds of billions (even trillions) of parameters, there may be sub-networks implementing various computations, and a wide variety of emergent behaviors corresponding to those computations. A similar situation occurs in research focused on understanding the human brain, where sub-networks of neuronal cells have been shown to localize specific capabilities. In cognitive science, researchers study such aspects of cognition without fully understanding the underlying neural circuitry, often modeling behavior without observing brain activity. Analogously, we seek to understand the structure of behaviors in large language models in the wild, without a full mechanistic understanding of their circuitry.
Bayesian inference is a key methodological tool of cognitive modeling, and recent work has framed in-context learning as Bayesian inference over models (Xie et al., 2022; Li et al., 2023b; Hahn & Goyal, 2023). Specifically, the posterior predictive distribution $p(y|x)$ in these works describes how an LLM produces output tokens $y$, given the context tokens, or prompt, $x$. The key assumption is that a context $x$ will activate latent concepts $c$ within a model according to their posterior probability $p(c|x)$, which the model marginalizes over to produce the next token $y$ by sampling from the posterior predictive distribution: $p(y|x)=\int_{c\in\mathcal{C}}p(y|c)\ p(c|x)$. This Bayesian model selection takes place in network activation dynamics, without changing model weights.
Following the framing above, in our experiments we assume a hypothesis space $\mathcal{H}$ that approximates the latent space of LLM concepts used when predicting the next token, i.e.,
$p(y|x)=\sum_{h\in\mathcal{H}}p(y|h)\ p(h|x),$
where $\sum_{h}$ can be changed to $\max_{h}$ to represent deterministic greedy decoding with an LLM temperature parameter of 0. If we assume that an LLM operates over next-token prediction only, then each generated token $y_{t-1}$ will be appended to the context when the following token is sampled, i.e. $y_{t}\sim p(y_{t}|x,y_{t-1})$. We specifically focus on Bernoulli processes, regular languages, Markov chains, and a simple memory-constrained probabilistic model as candidates for the hypothesis space $\mathcal{H}$ for estimating LLM concepts in random binary sequences. These candidates together capture a variety of behavioral patterns with varying complexity: Bernoulli processes serve as a baseline for a true random coin flip, where all elements in a sequence are independent and identically distributed; Markov chains are a minimal model capable of representing serial correlation patterns; the Gambler’s Fallacy bias in cognitive science (described below) emerges from a simple memory-limited model that draws the sequence towards a specified probability dependent based on previous samples; and Regular languages are well studied in theoretical computer science (Sipser, 1996; Chomsky, 1956), representing a simple class of non-random programs with known mechanisms of finite state automata (or, equivalently, finite state Markov processes). We use a subset of regular languages $(x)^{n}$, where $(x)$ is a short sequence of values, e.g., $(\texttt{010})^{n}$, where 00 maps to Heads and $1$ to Tails.
Learning Dynamics in Model Selection
Cognitive scientists have used Bayesian predictive and posterior distributions to model learning dynamics in human adults and children, as well as in non-human animals (Tenenbaum, 1998; Goodman et al., 2008; Ullman & Tenenbaum, 2020). A discrete hypothesis space in a probabilistic model can give clear and meaningful explanations to learning patterns analogous to mode collapse and phase transitions in deep learning. When behavior suddenly shifts from one pattern, or mode, of behavior to another, this can be understood as one hypothesis coming to dominate the posterior $p(h|x)$ as $x$ grows in scale (Ullman et al., 2012). For example, children between the ages of 3.5—5 years old learning to count undergo a dramatic conceptual shift from knowing the meanings of only a few number words (“one”, “two”) to a full inductive understanding of counting, which can be modeled as Bayesian model selection with a simplicity prior over models (Piantadosi et al., 2012). Such cognitive analysis of behavioral shifts as generated from a shift in posterior probability between one hypothesis to another parallels recent work on learning dynamics in training and prompting LLMs. E.g., sharp phase changes have been observed when training neural networks, corresponding to the formation of identifiable mechanisms such as modular addition circuits or induction heads (Nanda et al., 2023; Olsson et al., 2022; Zhong et al., 2023; Bai et al., 2023). ICL research has explored how few-shot prompting can significantly boost LLM performance in various domains, and how the particular exemplars provided in context determine its overall effectiveness (Brown et al., 2020; Zhao et al., 2020; 2021; Min et al., 2022; Lu et al., 2021; Zhou et al., 2022; Turpin et al., 2023). Work on chain-of-thought reasoning in LLMs demonstrates how a few exemplars of detailed solutions or even a simple prompt like “let’s think this through step-by-step” can dramatically impact model performance (Wei et al., 2022b; Kojima et al., 2022; Srivastava et al., 2022). For further discussion of related work, see Appendix A.
3 Algorithmic and Subjective Randomness
Next, we describe the domain of randomness that we explore in this work, and provide relevant background from computational cognitive science. To build intuition, consider: which sequence of coin flips is more random, $\mathtt{TTHTHHTH}$ or $\mathtt{HTTTTTTT}$? While the odds of a fair coin producing either sequence are the same (1 in 256), most people say that the first sequence “looks more random”. In a bias termed the Gambler’s Fallacy, people reliably perceive binary sequences with long streaks of one value as less random, and judge binary sequences with higher-than-chance alternation rates as being “more random” than truly random sequences (Falk & Konold, 1997). In addition to not using simple objective measures when judging the randomness of sequences, people also have a hard time generating sequences that are truly random  (Oskarsson et al., 2009; Griffiths et al., 2018; Gronchi & Sloman, 2021; Meyniel et al., 2016; Planton et al., 2021).
Figure 3: Determining whether a sequence is random can be viewed as search for a simple program that could generate that sequence. $\mathtt{HTTTTTTT}$ can be described with a short deterministic program $\mathtt{simpleSequence}$, and has a higher $p(h)$ according to a simplicity prior, compared to $\mathtt{TTHTHHTH}$ and $\mathtt{complexSequence}$. Both sequences can be generated by $\mathtt{randomSequence}$, but with lower likelihood $p(x|h)$.
Cognitive scientists studying Subjective Randomness model how people perceive randomness, or generate data that is subjectively random but algorithmically pseudo-random. One way to study this is to ask people whether a given data sequence was more likely to be generated by a Random process or a Non-Random process (Fig. 3). While the posterior distribution of all non-random processes is infinite, highly irregular, and includes every possible computable function, the problem of estimating this distribution can be simplified to finding the single most probable algorithm to approximate the full hypothesis space. If the hypotheses are programs which generate the data, a natural prior $p(h)$ is to assign higher probabilities to programs with shorter description lengths, or lower complexity. For example (Fig. 3), $\mathtt{HTTTTTTT}$ might be described with a shorter program—one H followed by sever Ts – than $\mathtt{HTTHTHHT}$, which cannot be compressed—H followed by two Ts, then one H, then one T, then two Hs, then one T. This optimization problem, searching for the shortest program consistent with data, is equivalent to computing the Kolmogorov complexity of a sequence $K(x)$ (Li et al., 1997) and has motivated the use of “simplicity priors” in a number of domains in computational cognitive science Chater & Vitányi (2003); Goodman et al. (2008); Piantadosi et al. (2012).
Following previous work (Griffiths & Tenenbaum, 2003; Griffiths et al., 2018), here we define subjective randomness of a sequence as the ratio of likelihood of that sequence under a random versus non-random model, i.e., $\mathtt{randomness}(x)=\log P(x|\mathtt{random})\ -\ \log P(x|\mathtt{non}\text{-}\mathtt{random})$. The likelihood given a truly random Bernoulli process will be $p(x|\mathtt{random})=2^{-|x|}$, since each of $2^{|x|}$ binary sequences of length $|x|$ has equal probability. The non-random likelihood $p(x|\mathtt{non}\text{-}\mathtt{random})=2^{-K(x)}$ denotes the probability of the minimal description length program that generates $x$, equivalent to Bayesian model selection: $P(x|\mathtt{non}\text{-}\mathtt{random})=\max_{h\in\mathcal{H}}p(x|h)\ p(h)$. In theory, the space $\mathcal{H}$ of non-random models encompasses every possible computation; in this work, we study a small subset of this space that includes formal languages and probabilistic models inspired by psychological models of human concept learning and subjective randomness (Griffiths et al., 2018; Nickerson, 2002; Hahn & Warren, 2009). For further background, see Appendix H.
4 Experiments
Figure 4: Prompt templates and GPT-3.5 Randomness Judgments   (Left) Prompt templates used for the Randomness Generation (Top) and Judgment (Bottom) tasks. $\{\mathtt{sequence}\}$ is substituted with a list of context flips $x$, e.g. $\mathtt{Heads,\,Tails,\,Tails,\,Heads}$. $\{\mathtt{p}\}$ is substituted with the probability of Tails, and $\{\mathtt{1-p}\}$ with the probability of Heads. (Right) In Randomness Judgment tasks, we observe sharp transitions in the predictive distribution $p(y=\mathtt{random}|x)$ for GPT-3.5, from high-confidence in $x$ being generated by a random process, to high confidence in a non-random algorithm, as additional data $x$ is provided from the same concept.
Randomness Generation and Judgment Tasks
[…]
In the Judgment task, $x$ includes the prompt question and the full sequence of flips. We systematically vary prompt context $x$ by varying the number of flips, denoted $|x|$. We test few-shot learning in LLMs by evaluating output behavior on specific bit sequences, for example $(\texttt{01})^{n}$ with varying $n$ (e.g. “$\mathtt{Heads,\,Tails,\,Heads,\,Tails,\,\dots}$”), as well as zero-shot language generation dynamics when $x$ is empty and $|x|=0$ (in practice, we initialize $x$ with a single flip to help the LLM match the correct format). Unlike typical cases of few-shot learning, where high-variance samples (‘shots’) are drawn from a broad category such as “chain-of-thought solutions” (Wei et al., 2022b; Brown et al., 2020), all data in a sequence $x$ corresponds to a single unobserved program.
Subjective Randomness Models
We compare LLM-generated sequences to a ground truth “random” Bernoulli distribution with the same mean ($\mu=\overline{y}_{\text{LLM}}$), to a simple memory-constrained probabilistic model, and to Markov chains fit to model-generated data $y$. The Gambler’s Fallacy is the human bias to avoid long runs of a single value (‘Heads’ or ‘Tails’) consecutively in a row (Falk & Konold, 1997; Nickerson, 2002). Hahn & Warren (2009) theorize that the Gambler’s Fallacy emerges as a consequence of human memory limitations, where ‘seeming biases reflect the subjective experience of a finite data stream for an agent with a limited short-term memory capacity’. We formalize this as a simple Window Average model, which tends towards a specific probability $p$ as a function of the last $w$ flips: $p(y|x)=\max(0,\min(1,2p-\overline{x}_{t-w\ldots t}))$
Sub-Sequence Memorization and Complexity Metrics
Bender et al. (2021) raise the question of whether LLMs are ‘stochastic parrots’ that simply copy data from the training set. We test the extent to which LLMs generate pseudo-random sequences by arranging memorized patterns of sub-sequences, rather than according to a more complex algorithm as our Window Average model represents. To measure memorization, we look at the distribution of unique sub-sequences in $y$. If an LLM is repeating common patterns across outputs, potentially memorized from the training data, this should be apparent in the distribution over length K sub-sequences. Since there are deep theoretical connections between complexity and randomness (Chaitin, 1977; 1990), we also consider the complexity of GPT-produced sequences. Compression is a metric of information content, and thus of redundancy over irreducible complexity (MacKay, 2003; Delétang et al., 2023), and neural language models have been shown to prefer generating low complexity sequences (Goldblum et al., 2023). As approximations of sequence complexity, we evaluate the distribution of Gzip-compressed file sizes (Jiang et al., 2022) and inter-sequence Levenshtein distances (Levenshtein et al., 1966).
Refer to captionRefer to captionRefer to caption
Figure 5: GPT-3.5 generates pseudo-random binary sequences that deviate from a Bernoulli process. (Left) Empirical conditional probabilities for a third-order Markov Chain fit to sequences $y$ generated by GPT-3.5 text-davinci-003, a Bernoulli process centered at the mean of GPT sequence $\overline{y}$, and our Window Average model ($w=5$). In the simulated Bernoulli process, edges are fairly uniform; the conditional probabilities for GPT-3.5 and the Window Average model demonstrate a similar non-uniform bias. (Right) Running averages for flip sequences generated by each model, where 00 denotes ‘$\mathtt{Heads}$’ and $1$ denotes ‘$\mathtt{Tails}$’. Compared to a Bernoulli process (top), sequences generating using GPT (middle) and those of our Window Average model (bottom) stay closer to the mean, repeating the same patterns more often.
Formal Language Learning Metrics
In our Formal Language Learning analysis, $x$ is a subset of regular expression repetitions of short token sequences such as $x\in(\texttt{011})^{n}$, where longer sequences $x$ correspond to larger $n$. This enables us to systematically investigate in-context learning of formal languages, as $|x|$ corresponds to the amount of data for inducing the correct program (e.g. $(\texttt{011})^{n}$) from Bernoulli processes, as well as the space of all possible algorithms. A Markov chain of sufficiently high order $k$ can implicitly represent a regular language in its weights. However, the number of parameters required is exponential in Markov order $k$, and weights are difficult to interpret. For example, a 3-order Markov chain fit to data $x=(\texttt{011})^{n}$ will represent $(\texttt{011})^{n}$ corresponds to high weights in the conditional probability distribution for $p(\texttt{0}|\texttt{11})$, $p(\texttt{1}|\texttt{01})$, $p(\texttt{1}|\texttt{10})$, and low values for other conditional probabilities. In randomness Judgment tasks, we assess formal concept learning by the dynamics of $p(y=\mathtt{random}|x=C^{|x|})$ as a function of $|x|$. In randomness Generation tasks, we asses concept learning according to the language model predictive distribution $p(y|x)$ over output sequences, inferred from next-token generation data: $p(y_{0\ldots T}|x)=p(y_{0}|x)\prod_{t}^{T}p(y_{t}|y_{0,\ldots t-1},x)$. Given a space of possible outputs $y$ with length $d$, $y\in\{0,1\}^{d}$, we estimate $p(y|x)$ by enumerating all $y$ up to some depth $d$, and computing $\hat{p}(y_{d}|x,y_{1,\ldots,d-1})=\frac{1}{N}\sum_{i}^{N}(y_{d}==1)^{(i)}$ as the fraction of $N$ responses that are “Tails”. We use token-wise log probabilities $p(y_{t}|y_{0\ldots t-1})$ from the OpenAI API where available for cost efficiency and since this is equivalent to repeated token samples. The probability table for $p(y_{t}|y_{0,\ldots t-1})$ is a weighted binary tree with depth $d$, where edges represent the next-token probability $p(y_{t}|y_{t-1})$, and paths represent the probability of a sequence of tokens. We estimate the predictive probability $p(y_{t}\in C|x,y_{0\ldots,t-1})$ assigned to a given regular language by computing the total probability mass for all paths in $y_{0\ldots d}$ that exactly match $C$. For example, with $C=(\texttt{011})^{n}$, there will be 3 paths $y_{0\ldots d}$ that exactly match $C$, out of $2^{d}$ possible unique paths.
5 Results
5.1 Subjectively Random Sequence Generation
Figure 6: GPT-3.5 shows a Gambler’s fallacy bias of avoiding long runs. (Top) Distribution of mean values of flip sequences ($\mu=\frac{1}{T}\sum_{t}y_{t}$) generated by GPT-3.5 (text-davinci-003) with the specified $p(Tails)$, compared with a Bernoulli process and our Window Average model with the same mean as the GPT-3.5 flips. Flips generated by GPT approximately follow the expected mean $p(Tails)$, but have lower variance than a Bernoulli distribution. (Bottom) Length of the longest run for each sequence, where a run is a sub-sequence of the same value repeating. In this case, we see a clear bias in GPT-3.5 to avoid long runs, with a similar pattern across all values of $p(Tails)$ despite the x-axis changing in scale.
In the GPT3.5 and higher (‘InstructGPT’) models — text-davinci-003, ChatGPT (gpt-3.5-turbo, gpt-3.5-turbo-instruct) and GPT-4 — we find an emergent behavior of generating seemingly random binary sequences (Fig. 5). This behavior is controllable, where different $p(\mathtt{Tails})$ values leads to different means of generated sequences $\overline{y}$. However, the distribution of sequence means, as well as the distribution of the length of the longest runs for each sequence, deviate significantly from a Bernoulli distribution centered at $\overline{y}$, analogous to the Gambler’s Fallacy bias in humans (see Fig. 6). Our Window Average model with a window size of $w=5$ partly explains both biases, matching GPT-generated sequences more closely than a Bernoulli distribution.
[…]
5.2 Sub-Sequence Memorization and Complexity
Figure 7: GPT repeats specific sub-sequences more often than chance. We find that GPT-3.5 repeats certain sub-sequences more often than is predicted by our Window Average model. While the Window Average model (green) generates fewer unique sub-sequences than a Bernoulli process (red), this does not account for the bias in GPT-3.5 (text-davinci-003, in blue) to repeat many of the same sub-sequences. This disparity increases with longer sub-sequences (see Fig. 20 in Appendix).
We find significant differences between the distributions of sub-sequences for GPT-3.5 -generated sequences and sequences sampled from a Bernoulli distribution(see Fig. 7). This difference is partly accounted for with a Window Average model with a window size $w=5$, although GPT repeats certain longer sub-sequences, for example length-20 sub-sequences, that are far longer than 5. However, the majority of sub-sequences have very low frequency, and though further experiments would be required to conclude that all sub-sequences are not memorized from training data, it seems unlikely that these were in the training set, since we find thousands of unique length-k (with varying k) sub-sequences generated at various values of $P(\mathtt{Tails})$. This indicates that GPT-3.5 combines dynamic, subjectively random sequence generation with distribution-matched memorization. We also note that with a specification of $P(Tails)=50\%$, but not $49\%$, $51\%$ or other values, sequences $y$ generated by GPT-3.5+ are dominated by repeating ‘Heads, Tails, Heads, Tails, …’. This pattern is consistent across variations of the prompts listed in Fig. 4, including specifying ‘fair’ or ‘unweighted’ instead of a ‘weighted coin’, and produces a visible kink in many cross-$p(\mathtt{Tails})$ metrics (Fig. 232219). For this reason, in Fig. 6 we show results for $P(Tails)=51\%$.
Across three metrics of sequence complexity — number unique sub-sequences, Gzip file size, and inter-sequence Levenshtein distance (see Fig. 2322 in Appendix) — we find that GPT-3.5+ models, with the exception of ChatGPT, generate low complexity sequences, showing that structure is repeated across sequences and supporting Goldblum et al. (2023); Delétang et al. (2023). By the metrics of mean Levenshtein distance and number of unique sub-sequences, ChatGPT generates higher complexity sequences than chance. We speculate that this phenomenon might explained by a cognitive model that avoids sampling with replacement.
Figure 8: With enough context, GPT-3.5 learns simple formal language concepts, transitioning from generating pseudo-random numbers to only generating values from the concept (Left) GPT-3.5 next-token predictive distribution $p(y|x)$ visualized as a binary tree, where red arrows correspond to Heads, blue arrows to Tails, and nodes matching the target concept $C=(011)^{n}$ are green. $p(y|x)$ changes with varying $|x|$, here $|x|=39$ and next-token predictions strongly follow the concept $C$. Also see Fig 1112 in Appendix. (Right) In-context learning dynamics for simple formal languages $x=(\texttt{HTH})^{n}$ (010) and $x=(\texttt{HTT})^{n}$ (011) as a function of context length $n=|x|$. Prediction accuracy computed as the total probability mass assigned to valid continuations of the formal language $x$, as a function of prediction depth $d=|y|$ and context length $|x|$. Curves shown are for $d=6$, where only 3 out of 64 total paths $y$ match concept $C$. Solid lines correspond to text-davinci-003 and dashed lines to gpt-3.5-turbo-instruct; note that learning curves for 010 and 011 flip between the two models. Also see Fig. 910 in Appendix.
5.3 Distinguishing Formal Languages from Randomness
GPT-3.5 sharply transitions between behavioral patterns, from generating pseudo-random values to generating non-random sequences that perfectly match the formal language (Fig. 8). We observe a consistent pattern of formal language learning in GPT-3.5 generating random sequences where predictions $p(y|x)$ of depth $d\geq 4$ are initially random with small $|x|$, and have low $p(y\in C|x)$ where $C$ is a given concept. This follows whether the prompt describes the process as samples from “a weighted coin” or “a non-random-algorithm”. We also find sharp phase changes in GPT-3.5 behavioral patterns in Randomness Judgment tasks across 9 binary concepts (Fig. 4). These follow a stable pattern of being highly confident in that the sequence is Random (high $p(y=\mathtt{random}|x$) when $x$ is low, up to some threshold of context at which point it rapidly transitions to being highly confident in the process being non-random. Transition points vary between concepts, but the pattern is similar across concepts (also see Fig. 1314 in Appendix).
6 Conclusion
We find strong patterns of random number Generation and Judgment in LLMs, including an emergent ability to generate low-variance pseudo-random binary sequences that can be partly explained with a memory-constrained moving window model, and partly as memorization and compression. We also find striking behavioral phase changes when GPT-3.5 learns to distinguish simple formal language concepts from randomness. This supports theories of ICL as Bayesian model selection and as under-specified program induction, and demonstrates how LLM behavioral patterns can change non-linearly with ICL as a function of context data. Future Cognitive Interpretability work may characterize concept learning dynamics underlying other LLM capabilities, including those with applications in usability and safety. If ICL is indeed analogous to hypothesis search in human cognition (or, equivalently, under-specified program induction) where new concepts can be learned in-context or computed in text generation trajectories, with no changes to model weights, a compelling research direction would be to further understand the features and limits of in-context learning dynamics using theories and tools from computational cognitive science.
Appendix A Related Work
Formal languages and transformers. A number of recent works explore how transformers and other neural language models learng formal languages (Delétang et al., 2022; Weiss et al., 2021; Shi et al., 2022; Allen-Zhu & Li, 2023; Bhattamishra et al., 2020; Wen et al., 2023; Liu et al., 2023; 2022a; Merrill & Sabharwal, 2023; Merrill et al., 2023; Merrill & Sabharwal, 2022). One common theme is that neural networks often learn ‘shortcuts’, degenerate representations of formal languages that fail out-of-samples.
In-Context Learning as Bayesian inference  A number of recent works frame ICL as Bayesian model selection (Xie et al., 2022; Li et al., 2023b; Akyürek et al., 2022; Hahn & Goyal, 2023; Bai et al., 2023). Two key differences in our work are: first, we analyze state-of-the-art LLMs based on behaviors alone, whereas prior work trains models from scratch on synthetic data and analyzes model parameters directly. Second, we consider Bayesian inference as an empirical modeling framework, as well as a theory, whereas these works only do the latter.
Mechanistic interpretability of transformer models  Prior work has characterized specific circuit-level implementations of simple high-level behaviors such as sequence copying, modular addition, and other primitive computational operations (Nanda et al., 2023; Goh et al., 2021; Geva et al., 2020; Belinkov, 2022; Li et al., 2022; Wang et al., 2022; Chughtai et al., 2023; Gurnee et al., 2023; Foote et al., 2023; Lubana et al., 2023; Lieberum et al., 2023; Barak et al., 2022; Liu et al., 2022b; Zhong et al., 2023). Our work differs in that we model hypothetical algorithms to characterize LM output behavioral patterns, without observing underlying activation patterns. We see this as analogous to cognitive science complementing neuroscience in the understanding of human cognition. We characterize the high-level “cognitive” representations in LLMs as a step towards connecting low-level explanations of neural circuits, such as induction heads, with sophisticated high-level behaviors that are characteristic of LLMs.
Language model evaluations  Our work resembles evaluation benchmarks such as BIG-Bench (Srivastava et al., 2022) that use behavior alone to evaluate LM understanding and reasoning. However, as described later, the domain of subjective randomness is fundamentally different in that there is no “correct” answer. Linguistic probing attempts to characterize the structure of LM representations, but unlike our work, is a function of hidden unit activations rather than output behavior.
LLM Text Generation Dynamics  Work on chain-of-thought reasoning in LLMs demonstrates how a few exemplars of detailed solutions or even a simple prompt like “let’s think this through step-by-step” can dramatically impact model performance (Wei et al., 2022b; Kojima et al., 2022; Srivastava et al., 2022), but typically only the model’s final answer is analyzed, not the trajectory of its intermediate steps. Our memory-constrained Window Average model, inspired by Hahn & Warren (2009), is similar in spirit to the claim of Prystawski & Goodman (2023), that ‘[chain-of-thought] reasoning emerges from the locality of experience’. Zhang et al. (2023a) demonstrate that invalid reasoning can snowball in LLMs, where hallucinations during intermediate steps lead to hallucinations in the final answer.
Random number generation in LLMsRenda et al. (2023) explore random number generation in LLMs, in addition to cursory explorations by (janus, 2022; Karpathy, 2023). These investigations do not analyze dynamics of sequence generation, nor do they ground their analysis, as we do, in theories of ICL as Bayesian model selection and the cognitive science of subjective randomness. Ortega et al. (2019) uses a similar domain as ours with random binary sequences and has a similar binary tree visualization over possible sequences, but they train models from scratch and analyze model hidden states, rather than behavioral trajectories as we do.
Bayesian program learning in cognitive science  Our work is inspired by computational cognitive science work that theoretically treats concepts as programs, and empirically uses structured Bayesian models to understand human cognition in various domains (Tenenbaum, 1998; Tenenbaum et al., 2011; Ullman et al., 2012; Ullman & Tenenbaum, 2020). We use models based on the cognitive science of subjective randomness (Falk & Konold, 1997), drawing particularly on the Bayesian program induction definitions of subjective randomness in Griffiths & Tenenbaum (2003; 2004); Griffiths et al. (2018). Our method of studying learning as probabilistic inference over formal languages with varying $|x|$ is also similar to Goodman et al. (2008); Piantadosi et al. (2012); Yang & Piantadosi (2022); Bigelow & Piantadosi (2016), who use more sophisticated grammar-based models of concept learning.
Appendix B Additional Experimental Details
All calls were made with the OpenAI API, using default parameters including — important to our analysis — a temperature parameter of $1.0$.
Since ChatGPT (not including gpt-3.5-turbo-instruct) and GPT-4 use a ChatCompletions API instead of Completions, we re-formatted the prompts in Fig. 4 to follow user/assistant/system prompt format. The following prompts were used:
System
Your responses will only consist of comma-separated "Heads" and "Tails" samples.
Do not repeat the userś messages in your responses.
User
[…]
Assistant
[…]
Appendix C Formal Language Learning, by Depth
Figure 9: Predictive distributions $p(y|x)$ by each LLM for Concept $(\texttt{010})^{n}$, at each prediction depth $d$. Colors correspond to different prediction depths, also refer to Figure 8. Note: text-ada-001 results are not shown since results did not follow the required format (‘Heads, Tails, …’) adequately to be analyzed.
Figure 10: Predictive distributions $p(y|x)$ by each LLM for Concept $(\texttt{011})^{n}$, at each prediction depth $d$. Colors correspond to different prediction depths, also refer to Figure 8.
Figure 11: Predictive distribution $p(y|x)$ trees with $d=6$ for concept $C=(011)^{n}$ with $|x|\in\{6,12,18\}$. Since $|x|$ is increasing by the same depth as the tree $\Delta_{|x|}=d=6$, the transition from generating pseudo-random numbers to deterministically repeating $011$ is visibly apparent. Also see Figure 8.
Figure 12: Predictive distribution $p(y|x)$ trees with $d=4$ for concept $C=(011)^{n}$ with $|x|\in\{1,9,18,24,39\}$. Models shown are text-davinci-002, text-davinci-003, and gpt-3.5-turbo-instruct. Also see Figure 8.
Appendix D Randomness Judgments
Figure 13: Randomness judgments across GPT models for 9 concepts Also see Fig. 4.
(Fig 13) text-davinci-003 shows a stable pattern of being highly confident (high token probability) in the process being random up to some amount of context $|x|$, at which point it rapidly transitions to being highly confident in the process being non-random, with transition points varying substantially between concepts. chat-gpt-3.5-instruct does not go through a stable high-confidence random period like text-davinci-003, and stable high-to-low confidence dynamics are observed for only a subset of concepts. The majority of earlier GPT models (text-davinci-002, text-davinci-001, text-curie-001, text-babbage-001) show no ’formal language learning’, at all. However, surprisingly OpenAI’s smallest available GPT model text-ada-001 shows S-shaped in-context learning dynamics, with the peak close to .5 instead of 1.0 as in text-davinci-003. Additionally, the learning dynamics and transition points for all concepts appear nearly identical, approximately at $|x|=50$, and some concepts show less stable "non-random" patterns for larger $|x|$.
Figure 14: Randomness Judgment $p(y=\mathtt{random}|x)$ dynamics for each concept tested, for text-davinci-003 and gpt-3.5-turbo-instruct
Appendix E Random Sequence Generation by GPT Model
Figure 15: 50 sequences sampled by each GPT model, for each $p(Tails)$. Color is assigned according to specified $p(Tails)$. Red dotted lines are drawn for each $p(Tails).$
Figure 16: 50 sequences sampled by text-davinci-003, for each $p(Tails)$, compared with samples from Bernoulli and Window Average models fit to $y_{LLM}$ for each $p(Tails)$. Color is assigned according to specified $p(Tails)$.
Figure 17: Probability $p(\mathtt{Tails})$ bias across LLMs text-davinci-003 and GPT-4 models are least biased relative to the specified $p(\mathtt{Tails})$ (x-axis). In the left figure, error bars represent the maximum and minimum sequence means $\overline{y}$ for each $p(\mathtt{Tails})$.
Figure 17: Probability $p(\mathtt{Tails})$ bias across LLMs text-davinci-003 and GPT-4 models are least biased relative to the specified $p(\mathtt{Tails})$ (x-axis). In the left figure, error bars represent the maximum and minimum sequence means $\overline{y}$ for each $p(\mathtt{Tails})$.
Fig. 15 and the left side of Fig. 17 demonstrate that text-davinci-003 and GPT-4 models not only are more controllable, following the correct probability more closely on average, but also have substantially lower variance than ChatGPT, which is both less controllable and has more variability in its distribution of responses. Further, GPT-4 is lower variance than text-davinci-003, with sequences staying even closer to their means $\overline{y}$.
Appendix F Gambler’s Fallacy Metrics by GPT Model
Figure 18: Gambler’s Fallacy histograms for ChatGPT (Top) and GPT-4 (Bottom). Also see Fig. 6.
ChatGPT shows no clear Gambler’s Fallacy bias, whereas GPT-4 does show this pattern, but is less pronounced than text-davinci-003 (Fig. 18).
Figure 19: Comparing metrics of Gambler’s Fallacy across probabilities and LLMs (Left) The mean longest run for each sequence $y$, at each specified probability $p(\mathtt{Tails})$, where a run is a consecutive sub-sequence of the same flip repeating multiple times in a row. (Right) The mean alternation rate for each LLM, where alternation rate is the fraction of consecutive flips that are not equal $p(y_{t}\neq y_{t-1})$.
Figure 19: Comparing metrics of Gambler’s Fallacy across probabilities and LLMs (Left) The mean longest run for each sequence $y$, at each specified probability $p(\mathtt{Tails})$, where a run is a consecutive sub-sequence of the same flip repeating multiple times in a row. (Right) The mean alternation rate for each LLM, where alternation rate is the fraction of consecutive flips that are not equal $p(y_{t}\neq y_{t-1})$.
In both plots of Fig. 19, we observe that text-davinci-003 shows a Gambler’s Fallacy bias across $p(\mathtt{Tails})$, of higher-than-chance alternation rates and shorter runs; ChatGPT (gpt-3.5-turbo-0613) produces more tails-biased and higher-variance sequences $y$ when $p(\mathtt{Tails})>50\%$; GPT-4 and gpt-3.5-turbo-instruct interpolate between the two distinct trends of text-davinci-003 and ChatGPT. The red dotted line represents a Bernoulli process with mean $p(\mathtt{Tails})$.
It is unclear how the capabilities we identify are implemented at a circuit level, or why they only seem to emerge in the most powerful and heavily tuned GPT models. For the latter, one hypothesis is that internet corpora contain text with human-generated or human-curating subjectively random binary sequences, and fine-tuning methods such as instruction fine-tuning, supervised fine-tuning, and RLHF make LLMs more controllable, enabling them to apply previously inaccessible capabilities in appropriate circumstances. Another hypothesis is that these fine-tuning methods bias LLMs towards non-repetitiveness, or induce some other general bias that plays a role in the in-context learning dynamics we observe in our particular domain. We hope that future work in cognitive and mechanistic interpretability will shed further light on these questions.
Appendix G Memorization, Compression, and Complexity
Figure 20: Distribution of unique sub-sequences for text-davinci-003 for varying sub-sequence lengths Also see Fig. 7.
Figure 20: Distribution of unique sub-sequences for text-davinci-003 for varying sub-sequence lengths Also see Fig. 7.
Figure 21: Distribution of unique sub-sequences for text-davinci-003, with additional models, varying sub-sequence lengths MC-2, MC-5, and MC-10 are Markov Chain models fit to GPT-3.5 flips, with orders $k=\{2,5,10\}$
Figure 21: Distribution of unique sub-sequences for text-davinci-003, with additional models, varying sub-sequence lengths MC-2, MC-5, and MC-10 are Markov Chain models fit to GPT-3.5 flips, with orders $k=\{2,5,10\}$
In Fig. 21, we show that Markov chains of high order $k$ can account for the sub-sequence distribution, but this only applies when $k<=w$ where $w$ is the sub-sequence length, and the Markov chains can effectively memorizing the sub-sequence distribution of $y$.
Figure 22: GPT-4 repeats the same sub-sequences more often than other GPT models (Left) Number of unique length-10 and length-20 sub-sequences as a function of specified probability $p(\mathtt{Tails})$, across all sequences $y$ (note: $|y|=50$) generated by each GPT model. (Right) The same distributions, with the y-axis normalized by dividing by the same metric (appended Gzip size, or mean Levenshtein distance) for a Bernoulli distribution centered at $\overline{y}$, to control for sequence compression being correlated with probability, e.g. with $\overline{y}=0.99$, the same sub-sequences of only ‘Tails’ flips will appear many times.
Across both unnormalized and normalized distributions of unique sub-sequences (Fig. 22), we find that GPT-4 repeats the same length-10 sub-sequences significantly more than the other models, and both ChatGPT-based models (gpt-3.5-turbo-0613, gpt-3.5-turbo-instruct) follow different patterns for $p(\mathtt{Tails})<50\%$ and $p(\mathtt{Tails})>50\%$, even when controlling for sequence bias (Right). The only model that generates more unique sub-sequences than chance (above dotted line) is ChatGPT (gpt-3.5-turbo-0613).
Figure 23: GPT-generated sequences have lower complexity than Bernoulli sequences
As a coarse approximation of sequence complexity, we use Gzip file size of appended sequences $gzip(y:y^{\prime}:y^{\prime\prime}:\ldots)$ and mean Levenshtein distance between sequences $d(y,y^{\prime})$. Gzip (Deutsch, 1996), a common algorithm for file compression that is highly optimized for compressing strings with redundancy into small file sizes, and Gzip file size has been found to be an effective feature extractor for NLP (Jiang et al., 2022). Levenshtein distance Levenshtein et al. (1966) is a measure of edit distance between two strings.
Since sequence compression is highly correlated with probability, e.g. all sequences with $\overline{y}=0.99$ will be highly compressible, we normalize the distribution of both plots in Fig. 23 by dividing by the same metric (appended Gzip size, or mean Levenshtein distance) for a Bernoulli distribution centered at $\overline{y}$. For all GPT models except ChatGPT, generated sequences have smaller Levenshtein distance than a Bernoulli process. This is evidence that these LLMs are using memorized sub-sequences (‘parroting’), since sequences have repeated structure. On the other hand, ChatGPT produces more dissimilar sequences than chance, suggesting higher complexity. In Gzip file size, however, we see a lower-complexity bias in all LLMs (except for a few higher values of $p(\mathtt{Tails})$), to varying degrees, produce data $Y$ that is more compressible than data from an equal probability Bernoulli process.
Appendix H Background on Algorithmic and Subjective Randomness
Randomness of a sequence $x$, defined in terms of Bayesian model comparison between the class of non-random models with the class of random models, can be translated to be the difference between the sequence length $|x|$ and the algorithmic complexity, or Kolmogorov complexity of the sequence $K(x)$.
[…]
$=K(x)-|x|$
The likelihood given a truly random Bernoulli process $p(x|\text{random})=2^{-|x|}$ since sequences of equal length have equal probability and there are $2^{|x|}$ binary sequences of length $|x|$. This can be thought of as a uniform prior over programs, where every program is an exact copy of the output string.
The likelihood of $x$ given the space of non-random processes marginalizes over the posterior of all non-random programs (hypotheses) $\mathcal{H}$:
$p(x|\text{non-random})=\sum_{h\in\mathcal{H}}p(h)\ p(x|h)$
A natural prior for programs $p(h)$ is the description length of that program, where common metrics used in software engineering such as lines of code or number of functions can be seen as practical estimations of program description length.
If we assume $p(x|h)$ is a binary likelihood, that is:
$p(x|h)=\begin{cases}1&\text{if }h\text{ generates }x\\ 0&\text{otherwise}\end{cases}$
and we simplify the problem to finding the maximum a-priori hypothesis $h$, and set a prior over hypotheses (programs) proportional to their length $p(h)=2^{-|x|}$, this equates to finding the program with lowest Kolmogorov complexity $K(x)$:
$P(x|\text{non-random})\approx\max_{h}p(h)\ p(x|h)=2^{-K(x)}$
where Kolmogorov complexity $K(x)$ is defined as the description length of the shortest program that generates $x$ as output:
$K(x)=\operatorname*{argmin}_{\{p\in\Sigma^{*}|Evaluate(p)=x\}}|p|$
The notation $p\in\Sigma^{*}$ is analogous to $h\in\mathcal{H}$, but refers to a formal alphabet $\Sigma$ that programs are comprised of. In the general case, Kolmogorov complexity $K(x)$ is uncomputable due to the halting problem, since the expression $Evaluate(p)=x$ might run forever if $p$ has an infinite loop.